How many grams of ammonium carbonate are needed to decompose in order to produce 6.52 g of carbon dioxide
1
Expert's answer
2019-03-01T05:01:20-0500

Solution:

The balanced equation for reaction of ammonium carbonate decomposition can be written as follows

(NH4)2CO3 (s) → 2NH3 (g) + CO2 (g) + H2O (g)

"\frac{m((NH_{4})_{2}CO_{3})}{M((NH_{4})_{2}CO_{3})}=\frac{m(CO_{2})}{M(CO_{2})}"

"m((NH_{4})_{2}CO_{3})=\frac{m(CO_{2})\cdot M((NH_{4})_{2}CO_{3})}{M(CO_{2})}"

"M((NH_{4})_{2}CO_{3})=2\cdot (1\cdot 14+4\cdot 1)+1\cdot12+3\cdot16=96\;amu"

"M(CO_{2})=1\cdot12+2\cdot16=44\;amu"

"m((NH_{4})_{2}CO_{3})=\frac{6.52\cdot 96}{44}\approx 14.2 \; g"


Answer:

14.2 g